# 2Meituan, Shanghai, China

### v1

@@ 3.3. EvalAgent and Evaluation
  • Unit Test: As in previous benchmark studies Jimenez et al. (2023), unit tests remain the most effective approach for verifying the functional correctness of individual components and modules within the code. PRDBench retains this type of testing and provides auxiliary pytest-based test scripts. The EvalAgent simply runs the pytest command to execute these tests.
@@
  • Shell Interaction: For tasks involving command-line interaction or external system operations, the EvalAgent executes predefined shell commands and compares the actual outputs to the expected results, ensuring the code correctly handles system-level operations and user inputs. PRDBench supplies comprehensive simulated user input files and program entry commands for this purpose.
@@
  • File Comparison: For project-level tasks that generate files or require specific directory structures, the EvalAgent compares the produced files against reference solutions, checking for correctness in content, format, and organization. PRDBench provides the raw data for target file generation, the corresponding Python commands, and the reference solution files.
By leveraging these tools, the EvalAgent can automatically adapt to specific code repositories and complete all required tests. All test categories and supporting files are provided within PRDBench, which greatly reduces the complexity of evaluation. EvalAgent only needs to execute each test and analyze the results. This automated evaluation process generates detailed reports on the performance and correctness of code agent submissions.
@@ 4. Experiments and Results
4.1. Data Statistics
PRDBench comprises 50 project-level tasks, encompassing a total of 1,262 scoring points. Specifically, there are 409 unit test points, 729 shell interaction points, and 124 file comparison points. Each PRD contains, on average, 105.22 lines. Figure 4(a) presents the distribution of scaffold line counts recorded during the annotation.
For all PRDs, we utilize GPT-4 to assign a domain label to each task. From these labeled tasks, we select 50 tasks that span 20 distinct domain labels to construct PRDBench. The distribution of these domain labels is illustrated in Figure 4(b). To ensure broad domain coverage, we include at least one representative task from less common domains. For more prevalent areas of Python development, such as data processing and machine learning, we select tasks that require a variety of technical approaches and sub-skills. This sampling strategy ensures that PRDBench provides comprehensive and balanced coverage across different Python application domains.
(a) Scaffold line counts distribution.
(b) Domain distribution of PRDBench.
Figure 4. Data statistics of PRDBench.
@@ 4.2. Code Agents
Minimal Basic code tools Qwen3-Coder (Team, 2025c)
GPT-5 (OpenAI, 2025b)
Claude-3.7-Sonnet (Anthropic, 2025a)
Gemini-2.5-pro (Team, 2025a)
Commercial Gemini CLI (Mullen and Salva, 2025) Gemini-2.5-pro (Team, 2025a)
Claude Code (Anthropic, 2025a) Claude-4/3.7 Hybrid (Anthropic, 2025b)
CodeX (OpenAI, 2025a) GPT-5 (OpenAI, 2025b)
Qwen Code (Team, 2025b) Qwen3-Coder (Team, 2025c)
EvalAgent EvalAgent Qwen3-Coder (Team, 2025c)
@@
    Minimal Agents [^2] Code is available in https://github.com/AGI-Eval-Official/Minimal-CodeAgent., implemented using the Agent Development Kit (ADK). These agents are equipped with essential tools for file manipulation, bash scripting, and Python execution. We integrate state-of-the-art LLMs (Claude-3.7-Sonnet, Gemini-2.5-Pro, Qwen3-Coder-480B-A35B, and OpenAI GPT-5) to systematically assess their core capabilities in strategy implementation. Hereafter we denote the models as Claude, Gemini, Qwen3-Coder and GPT-5 respectively for simplicity.
@@ 4.3. Experimental Setups
We select Qwen3-Coder Team (2025b) as the backbone LLM for EvalAgent. Qwen3-Coder is fully open-source, which facilitates future fine-tuning and further experiments for Agent-as-a-Judge scenarios within PRDBench. For code agents, we select the latest model from each family that is compatible with the agent framework. However, Claude-4-Sonnet occasionally encounters tool call issues that result in no code being generated, as frequently reported by the community. Therefore, we use Claude-3.7-Sonnet in minimal code agent. For all LLMs used in our work, we set temperature to be $0.1$, max token identical to their official APIs’ setting. We set top-p to $1.0$, Top-k to be $100$, and presence penalty to be default to the API.
@@
Each code agent is executed using a Python virtual environment that contains necessary and useful packages for the agents. If the code agent need other packages, we allow it use pip to install them.
4.4. Main Results
@@
Minimal Agent
GPT-5 55.81 60.15 4.34
Claude 45.50 49.07 3.57
Gemini 14.27 15.99 1.72
Qwen3-Coder 37.60 46.98 9.38
Commercial Agent
CodeX 56.23 50.24 -6.99
Claude Code 36.60 45.53 8.93
Gemini CLI 16.35 21.59 5.16
Qwen Code 39.59 35.69 -3.90
Table 3 presents the performance of all evaluated code agents on PRDBench. In Round 1, we provide the PRD and the relevant evaluation materials, allowing each code agent to implement the entire project based on the existing interfaces. This round assesses the development capability of the code agents. In Round 2, we provide the EvalAgent report from the first round along with the Round 1 code, enabling each code agent to analyze and revise the code according to the feedback and identified issues. This round evaluates the debugging capability of the code agents.
Table 4. Cost statistics of code agents. Code indicates the number of lines generated in Round 1 (development) and the number of lines modified in Round 2 (debug). Relevant log data for CodeX is unavailable, so its results are not reported.
Round 1 (DEVELOPMENT) Round 2 (DEBUG)
Time (s) Input Output Code (lines) Time (s) Input Output $\Delta$ Code (lines)
Minimal Agent
GPT-5 1517.75 1506252.18 43467.62 854.49 1165.25 1888893.43 32358.57 774.04
Claude 717.50 1775631.15 33398.27 1975.35 380.37 599139.68 22760.68 285.96
Gemini 760.54 609661.54 27374.89 667.78 1370.54 3419419.51 32752.20 1579.06
Qwen3-Coder 892.35 1136848.60 17209.19 982.67 495.48 1835992.38 14176.74 917.50
Commercial Agent
CodeX - - - 651.62 - - - 101.18
Claude Code 1309.84 4113143.56 46669.50 1751.80 729.74 2140884.40 19593.26 112.68
Gemini CLI 2740.70 834125.00 20238.96 362.13 2067.61 5776370.08 24611.50 546.12
Qwen Code 1183.42 3108353.62 31400.28 1183.42 9582.81 988041.65 5604.89 31.00
@@
Our main findings are as follows:
(1) The coding ability of the underlying LLM significantly impacts the development performance of the code agent. We observe that the relative ranking of minimal agents in round 1 is consistent with that of the commercial agents, indicating a strong correlation between the LLM’s capabilities and the agent’s development outcomes.
(2) Adapting the code agent framework to LLMs can further enhance development performance. While commercial agents typically outperform minimal agents based on the same LLM, there are reasonable exceptions. For example, since Claude-3.7-Sonnet does not appear to be specifically optimized for Claude Code, Claude Code achieves lower scores than the ADK-based Claude agent in our experiments. Additionally, Gemini-2.5-Pro and Qwen3-Coder demonstrate better generalization and more stable performance across different code agent frameworks.
(3) Debugging and initial development require distinct capabilities from code agents. In the Debug phase (Round 2), we provide the previous round’s test report to guide code agents in targeted fixes. Minimal code agents demonstrate more stable improvements during debugging, consistently increasing their accuracy after applying corrections. In contrast, CodeX experiences a dramatic drop in performance in the DEBUG phase due to interface inconsistencies introduced during modification, which led to additional errors. These observations suggest that effective debugging requires not only error identification but also careful maintenance of code structure and interfaces, highlighting the importance of robust error analysis and correction strategies in code agent designs.
4.5. Code Agent Analysis
@@ 4.5.1. Cost of code agents
Table 4 presents the resource consumption statistics for each code agent, including time spent, input and output token counts, and lines of code generated or modified during both development (Round 1) and debugging (Round 2). Overall, commercial agents tend to consume more time than minimal agents during both phases. For each backbone LLM, the resource consumption patterns of its minimal and commercial agent implementations are generally consistent. Notably, Gemini exhibits significantly lower input token consumption in the development phase compared to other models, whereas Qwen3-Coder consume the least output token. However, in the debugging phase, Gemini’s input token usage is more than twice that of other agents, indicating that the development and debugging phases pose distinct challenges for code agents in terms of resource demands. Regarding code modification, Claude Code and Qwen Code made relatively few changes during debugging, while GPT-5 and Gemini demonstrated much larger code modifications. This suggests that different code agents adopt varying strategies or exhibit different levels of responsiveness when addressing feedback and implementing code revisions. Some agents may focus on targeted, minimal fixes, while others opt for more extensive refactoring or rewriting in response to evaluation reports.
In summary, the observed differences in resource consumption and code modification patterns highlight that development and debugging are fundamentally different for code agents, requiring distinct approaches and optimization strategies. The results also suggest that agent architecture and underlying LLM characteristics strongly influence how agents allocate resources and perform code updates during the evaluation process.
@@ 4.5.2. Error Analysis of Code Agent
4.6. EvalAgent Analysis
4.6.1. Human Alignment
We assess the reliability of the evaluation agent through manual annotation. Specifically, we randomly select 16 code submissions, each corresponding to a distinct problem, and employ two annotators to score the outputs of the code agent according to the criteria scheme from PRDBench. These 16 codes are generated by different code agents and span different problems.
The results of human alignment are summarized in Table 5. In total, we examine 282 distinct test cases. Of these, 230 test cases (81.56%) achieve perfect score alignment between the annotators and the EvalAgent, indicating that EvalAgent’s evaluation results are generally reliable. For 9 cases (3.2%), the absolute difference in scores is 1, and for 43 cases (15.2%), the absolute difference is 2. We analyze the alignment rates across different projects and observe substantial variance (variance = 774.66), indicating significant differences in score alignment between projects. The standard deviation of alignment rates is also relatively high (27.83%), with values ranging from 0% to 100%. This suggests that the consistency between machine and human scoring can vary greatly depending on the specific project or model. The result implies that the stability of EvalAgent remains a major challenge, as an error occurs in one test case, subsequent evaluations within the same project may be affected due to contextual dependencies, potentially leading to distorted overall assessment for that problem. This phenomenon is also observed in other LLM-driven agents Lin et al. (2025), and attributed to propagation chains.
In addition, we examine the alignment rates among the three major test case types: file comparison (84.62%), shell interaction (82.55%), and unit test (79.44%). The alignment rates are comparable across these categories, indicating that the evaluation difficulty for EvalAgent is similar among the three types. This demonstrates that our three test formats, as well as the commonly used unit test benchmarks, are well-suited for reliable assessment by EvalAgent.
Table 5. Alignment between Human-as-a-Judge and Agent-as-a-Judge.
Total Unit Test Shell Inter. File Comp.
Human-as-a-Judge
Number 282 107 149 26
Alignment 81.56% 79.44% 82.55% 84.62%
4.6.2. Cost of EvalAgent
Table 6. Average cost of EvalAgent. The input and output are quantified by the number of tokens, and the API cost is estimated according to the official pricing of Qwen3-coder.
Time (s) Input Output API Cost(\$)
EvalAgent 425.62 124,2440 8,825 2.68
Table 6 shows the average time and token cost for EvalAgent. On average, EvalAgent generates a comprehensive testing report in just 7 minutes at an API cost of \$2.68 per problem, compared to 0.5 to 1 hour for human annotators (excluding writing reports). This demonstrates that EvalAgent and PRDBench greatly improve evaluation efficiency while keeping costs low.
4.7. Free Development

### v2

@@ 3.3. Data Statistics
  • Unit Test (408 metrics): As in previous benchmark studies jimenez2023swe, unit tests remain the most effective approach for verifying the functional correctness of individual components and modules. PRDBench retains this rigorous testing paradigm and provides auxiliary pytest-based test scripts.
@@
  • Shell Interaction (732 metrics): For tasks involving command-line interaction or external system operations, predefined shell commands are executed to compare actual outputs against expected results. This ensures the code correctly handles system-level operations and user inputs. PRDBench supplies comprehensive simulated user input files and program entry commands for this purpose.
@@
  • File Comparison (118 metrics): For project-level tasks that generate files or require specific directory structures, the produced artifacts are compared against reference solutions to verify correctness in content, format, and organization. PRDBench provides the raw data for target file generation, the corresponding Python commands, and the reference solution files.
4. PRDJudge
In this section, we present PRDJudge, a specialized evaluation agent developed to automatically and reliably assess code submissions on PRDBench. We elaborate on its tool-augmented evaluation framework, and the fine-tuning methodology designed to align its scoring with human standards.
4.1. Evaluation Framework
Figure 5. Overview of PRDJudge. PRDJudge executes tests based on the criteria scheme using various tools, compares outputs (files or results) with expected outputs, and generates a report for the submitted code.
To automate the validation of code generated by code agents, we introduce PRDJudge, a lightweight evaluation agent tailored for the predefined criteria schemes of PRDBench (illustrated in Figure 5). For each item in the criteria scheme, PRDJudge leverages the executable terminal commands and expected outputs provided by the benchmark to conduct systematic verification.
To interact effectively with the testing environment, PRDJudge is equipped with six core tools, including file reading and writing, command-line execution, image handling, and a specialized judge tool. Among these, the deal graph tool offers a multimodal LLM interface (powered by GPT-4o in this paper), allowing the agent to process visual outputs by submitting both the image and a descriptive prompt to obtain verification results—this is particularly useful for tasks requiring UI or plot validation. Furthermore, the judge tool is specifically designed for PRDBench. It accepts files containing simulated user inputs, enabling the generation of comprehensive terminal logs for thorough analysis without the need for manual input specification.
By leveraging this toolset, PRDJudge automatically adapts to specific code repositories and seamlessly executes the three categories of tests (Unit Test, Shell Interaction, and File Comparison) defined in Section 3.3. Since all supporting files and test categories are self-contained within PRDBench, the complexity of evaluation is greatly reduced. PRDJudge only needs to execute each test, analyze the execution logs or file differences, and generate a detailed report on the performance and correctness of the code agent submissions. However, accurately interpreting complex execution logs and aligning the final scores with human quality assurance standards requires exceptional reasoning capabilities, which motivates the training of a specialized evaluation model, as detailed in the following section.
4.2. Fine-tuning PRDJudge
4.2.1. Human Annotation Protocol
To establish a high-quality ground truth for both training data filtering and final evaluation, we employ a rigorous human annotation process. Detailed information of our annotators is provided in Appendix A. Each scoring point in the benchmark is independently evaluated by two human annotators. In cases of disagreement, a third senior annotator serves as an arbitrator to make the final decision. The annotation follows a strict three-tier ordinal scoring system:
  • Score 2 (Pass): The code executes perfectly, and the output strictly matches the expected requirements specified in the criteria scheme.
  • Score 1 (Partial): The code runs successfully (i.e., the interfaces are correct and no execution crashes occur), but the final output or behavior does not fully align with the expected results.
  • Score 0 (Fail): The code fails to execute entirely, typically due to syntax errors, missing dependencies, or severe runtime exceptions.
4.2.2. Training Data Construction
The training data for PRDJudge consists of high-quality evaluation trajectories. To collect these trajectories, we deploy a highly capable model, Qwen3-Coder-480B-A3B qwen3technicalreport, as the initial judge within the PRDJudge framework. We collect evaluation trajectories from the outputs of 11 different code agents previously tested on PRDBench. To ensure diversity in the training distribution, we randomly sample 8 repositories for each agent to form the training pool.
Crucially, to prevent the model from learning hallucinated or flawed reasoning, we apply a strict two-stage filtering mechanism. Initially, Qwen3-Coder-480B-A3B generated a total of 2,147 evaluation trajectories. In the first stage, we only retain the 1,742 trajectories where the final score predicted by the model exactly matches our human-annotated ground truth. It is worth noting that while the generation of PRDBench tasks is highly efficient and scalable (as discussed in Section 1), establishing this ground truth required a substantial human annotation effort. However, we consider this intensive labeling process a necessary, one-time investment to distill a reliable automated judge, which subsequently enables zero-cost, scalable evaluations for any future code agents. Furthermore, to ensure the efficiency and accuracy of the agent’s testing trajectories, we apply an additional rule-based filter to eliminate trajectories that are excessively long or exhibit invalid tool usage. Although these discarded trajectories achieved the correct final score, their inaccurate tool usage could introduce noise into the evaluation agent’s scoring behavior. Ultimately, this rigorous filtering process yields a final dataset of 911 high-quality training trajectories for fine-tuning.
@@ 5.1. Code Agents
Minimal Basic code tools Qwen3-Coder-480B-A3B (qwen3technicalreport)
GPT-5.2 gpt5.2
Claude-4.5-Sonnet anthropicclaude45
Gemini-3-pro gemini3
Kimi-K2-Instruct-0905 kimiteam2025kimik2openagentic
DeepSeek-V3.2
GLM-4.7 5team2025glm45agenticreasoningcoding
Minimax-M2 minimaxm2
Commercial Gemini CLI mullen2025geminicli Gemini-2.5-pro gemini25
Claude Code anthropic2025claude37 Claude-4.5-Sonnet anthropicclaude45
CodeX codex GPT-5 gpt5
Qwen Code qwencode Qwen3-Coder-480B-A3B qwen3technicalreport
EvalAgent PRDJudge Qwen3-Coder-30B-A3B qwen3technicalreport
@@
    Minimal Agents [^2] Code is available in https://github.com/AGI-Eval-Official/Minimal-CodeAgent., implemented using the Agent Development Kit (ADK). These agents are equipped with essential tools for file manipulation, bash scripting, and Python execution. We integrate state-of-the-art LLMs (Qwen3-Coder-480B-A3B, GPT-5.2, Claude-4.5-Sonnet, Gemini-3-Pro, Kimi-K2-Instruct-0905, DeepSeek-v3.2, GLM-4.7, and Minimax-M2) to systematically assess their core capabilities in strategy implementation. Hereafter we denote the models as Qwen3-Coder, GPT-5.2, Claude-4.5, Gemini-3-Pro, Kimi-K2, DeepSeek-v3.2, GLM-4.7, and Minimax-M2 respectively for simplicity.
@@ 5.2. Experimental Setups
For code agents, we select the latest model from each family that is compatible with the agent framework. For all LLMs used in our work, we set the temperature to be $0.1$, and the max tokens identical to their official APIs’ setting. We set top-p to $1.0$, top-k to $100$, and presence penalty to be default to the API. Each code agent is executed using a Python virtual environment that contains necessary and useful packages for the agents. If the code agent need other packages, we allow it use pip to install them.
For PRDJudge training, we adopt Qwen3-Coder-30B qwen3technicalreport as our base model, balancing strong coding capabilities with deployment efficiency. To optimize training costs while maintaining performance, we employ Low-Rank Adaptation (LoRA) for parameter-efficient fine-tuning. The LoRA rank is set to 16. The model is trained for 1 epoch using a learning rate of $2.0\times 10^{-4}$ and a warmup ratio of 0.03. The fine-tuning process is conducted on a cluster of 8 NVIDIA H800 GPUs.
5.3. PRDJudge Evaluation
5.3.1. Evaluation Setup
To evaluate the reliability and generalizability of PRDJudge, we construct two complementary test sets with full human annotation. (1) In-Domain Test Set. We collect code submissions generated by the same 11 code agents used in the training data construction phase. For each agent, we randomly sample 2 repositories (different to training data), yielding a total of 513 scoring metrics for human annotation. This test set is designed to assess PRDJudge’s alignment with human judgment under familiar agent and repository distributions. (2) Out-of-Domain Test Set. To evaluate the generalization capability of PRDJudge beyond the training distribution, we additionally construct an out-of-domain test set, where the code agents whose outputs are evaluated did not appear in the training data. This set comprises 26 repositories spanning diverse domains, resulting in a total of 605 scoring metrics. This setting more closely reflects real-world deployment scenarios, where PRDJudge may encounter previously unseen agent behaviors and implementation styles.
All annotations undergo a rigorous two-round labeling process to ensure quality and consistency. Specifically, each metric is independently annotated by two annotators. For cases where both annotators reach the same score, we randomly sample 10% of such instances for a third-round quality check by a dedicated inspector. For cases where the two annotators disagree, the inspector reviews both annotations and provides a final adjudicated score, taking into account the reasoning of both parties. This multi-stage annotation pipeline ensures high inter-annotator reliability and minimizes subjective bias in the ground-truth labels.
5.3.2. Evaluation Metric
To assess the reliability of PRDJudge against human standards, we employ the Human Alignment Rate (HAR) as our primary evaluation metric. HAR calculates the exact match accuracy between the scores predicted by the evaluation model and the human-annotated ground truth. It intuitively reflects the overall correctness and reliability of the judge.
Table 3. Performance comparison of PRDJudge and baseline models on the In-Domain and Out-of-Domain test sets. The accuracy is measured by the Human Alignment Rate (HAR, %). Avg. Token and Avg. Time (in seconds) denote the average context length and inference time required per metric evaluation.
Model In-Domain Out-of-Domain Avg. Token Avg. Time
Human 95.83 95.83 - ¿ 600
Qwen3-Coder (480B) 90.32 87.91 91,985 114.96
Claude-4.5 90.64 88.10 166,386 174.64
GPT-5.2 89.76 87.09 33,182 133.96
Qwen3-Coder (30B) 70.52 55.46 151,411 162.34
PRDJudge (ours) 91.75 92.69 103,712 107.85
5.3.3. Human Alignment
Table 3 presents the evaluation results against human annotations, leading to three key conclusions: (1) PRDJudge achieves state-of-the-art human alignment. PRDJudge outperforms all baseline models, including massive proprietary LLMs, and approaches the human inter-annotator agreement upper bound. The massive improvement over its base model (Qwen3-Coder 30B) highlights the necessity of specialized training for complex, long-context PRD-based evaluation. (2) PRDJudge exhibits exceptional out-of-domain generalizability. Zero-shot baseline models experience a noticeable performance drop on the out-of-domain test set, indicating this set is intrinsically harder (likely due to complex edge cases from unseen, stronger agents). Despite this challenge, PRDJudge maintains highly stable performance. This proves it genuinely internalizes the underlying evaluation criteria rather than merely overfitting to the training agents’ specific error patterns. (3) PRDJudge offers superior computational efficiency. Alongside the highest accuracy, PRDJudge records the lowest average inference time. Compared to its base model, it significantly reduces token consumption and inference time, suggesting that specialized training enables more concise, targeted reasoning. This makes PRDJudge a highly practical and scalable solution for real-world automated evaluation.
5.3.4. Stability Analysis
To evaluate scoring stability, we conduct three independent evaluation runs for each metric and measure the Unanimous Agreement Rate (UAR) and Pairwise Agreement Rate (PAR). Detailed results are provided in Appendix B. PRDJudge demonstrates the highest stability among all evaluated models, achieving a UAR of 94.19% and a PAR of 96.07%. The superior consistency indicates that our specialized fine-tuning helps PRDJudge internalize a deterministic evaluation mapping from PRDs to code, significantly mitigating the reasoning variance and generation randomness typically observed in general-purpose LLMs.
5.3.5. Case Study
To intuitively understand why PRDJudge achieves superior performance and efficiency, we trace and analyze the tool-use trajectories of different models during the evaluation process. The complete tool usage distribution and case study is provided in Appendix C. First, PRDJudge follows a logical explore-then-evaluate workflow, prioritizing repository exploration and file reading to understand code logic before making judgments. In contrast, GPT-5.2 exhibits a blind action pattern, characterized by excessive command execution and file writing with minimal code inspection. Second, PRDJudge demonstrates high precision. While its base model (Qwen3-30B) suffers from tool spamming due to convoluted reasoning, PRDJudge requires significantly fewer, more targeted calls than even Claude-4.5. This confirms that specialized training enables the model to optimally gather evidence rather than merely increasing tool-use frequency.
5.4. Code Agent Evaluation
5.4.1. Overall Results
Table 4 presents agent performances across two phases: Round 1 (DEV) assesses zero-shot project implementation based on PRDs, while Round 2 (DEBUG) evaluates iterative refinement using PRDJudge feedback. Our main findings are: (1) Foundation models dictate zero-shot development ceilings. Frontier models (e.g., Claude-4.5, GPT-5.2) consistently dominate the DEV phase, indicating a strong correlation between intrinsic LLM reasoning capacity and the ability to generate project-level scaffolds from scratch. (2) Commercial frameworks constrain initial generation but excel in iterative debugging. Minimal agents generally outperform their Commercial counterparts in DEV, suggesting that complex tool-use guardrails in commercial frameworks restrict zero-shot generation. Conversely, commercial agents demonstrate superior debugging capabilities, effectively utilizing feedback for targeted fixes. (3) Unconstrained models risk performance degradation during long-context debugging. High-performing Minimal agents (e.g., Claude-4.5, DeepSeek-V3.2) experience substantial drops in the DEBUG phase. Without robust agentic scaffolding to manage global states, attempting localized modifications frequently breaks previously functioning modules, underscoring the necessity of structured designs for safe code maintenance. (4) LLMs exhibit divergent generation and refinement profiles. Within Minimal agents, we observe three archetypes: (i) strong zero-shot generation but high variance in context maintenance, causing debugging regressions (Claude-4.5, DeepSeek-V3.2); (ii) limited initial generation but strong instruction-following for localized fixes, yielding significant improvements (Kimi-K2, Gemini-3-Pro); and (iii) consistent performance across both phases (GPT-5.2, Qwen3-Coder). This suggests future agent designs must adapt to the specific reasoning characteristics of their backbone models.
@@ 5.4.1. Overall Results
Minimal GPT-5.2 62.49 69.00 +6.51
Claude-4.5 69.19 56.40 -12.79
Gemini-3-Pro 22.76 27.28 +4.52
Qwen3-Coder 43.84 48.29 +4.45
Kimi-K2 20.52 36.17 +15.65
DeepSeek-V3.2 40.11 24.80 -15.31
GLM-4.7 38.39 35.33 -3.06
Minimax-M2 17.60 24.75 +7.15
Commercial CodeX 62.09 65.02 +2.97
Claude Code 56.65 70.25 +13.60
Gemini CLI 11.29 21.51 +10.22
Qwen Code 39.91 35.95 -3.96
@@ 5.4.2. Cost of Code Agents
Appendix D details the computational expenditures of the evaluated LLM-based code agents across two experimental iterations, revealing distinct operational patterns. First, the generation and maintenance profiles diverge significantly across agent configurations. The execution of minimal LLM agents is characterized by high-volume initial code generation, followed by high volatility during the debugging phase, where modifications vary from minor syntax corrections to extensive logic restructuring within the context window. Conversely, commercial LLM frameworks operate under highly constrained generation parameters. Their maintenance phase is defined by localized modifications, which restricts code alterations to a stable and minimal scope regardless of task complexity. Second, the underlying drivers of computational overhead differ fundamentally. For minimal agents, token consumption and inference latency scale directly with the volume of generated code. In contrast, commercial agents incur substantial token expenditures and latency that remain largely decoupled from the actual volume of code modification. This resource distribution reflects an architectural reliance on complex tool-use reasoning, continuous environment synchronization, and iterative state management rather than baseline zero-shot drafting speed.
@@ 5.4.3. Error Analysis of Code Agent
5.4.4. Free Development